# Behaviour Policy Optimization: Provably Lower Variance Return Estimates for Off-Policy Reinforcement Learning

### v1

@@ 3.3 Variance reduction
Figure 1: (a) Sampling densities. (b) Distribution of estimates over 100,000 repeated trials.
@@ 4 Off-policy TD($\lambda$)
The analysis can be found in Appendix A.4. This property is important, as with appropriate hyperparameter settings we can achieve unbiasedness. We treat $\bar{c}$ and $\bar{\rho}$ as hyperparameters; generally this result indicates that a good approach is to set $\bar{\rho}\geq\bar{c}\geq 1$. Additionally, $\lambda$ should not stray too far from 1 as it biases our estimate when we rely on value estimates. We continue with the the following result.
@@
Theorem 1 ensures that while we search over the space of policies $\Lambda$ our estimator remains unbiased (under certain conditions). The proof can be found in Appendix A.5. Now we provide a proof of variance reduction for $G_{t}^{\text{TIS}}$ under the behaviour policy $\hat{\mu}$ designed in (5).
@@
And, $\epsilon(s):=\gamma^{2}\mathbb{E}_{A_{t}\sim\hat{\mu}_{t}}\Big[\rho_{t}^{2}\mathbb{E}_{S_{t+1}\sim p}\big[\epsilon(S_{t+1})\big]\Big]$
The proof can be found in Appendix A.6. Noting here that $c(s)$ is always non-negative by Jensen’s inequality, which guarantees non-negativity of $\epsilon(s)$ and thus the property of variance reduction. $c(s)=0$ occurs only in the degenerate case when all actions have the same $\hat{q}_{\pi}$ for state $s$. Finally, it remains to show how $\hat{q}_{\pi}$ can be learnt,
@@
The proof here is a straightforward adaptation of (liu2024efficient) and intermediate results from the other theorems, c.f., Appendix A.7. Theorem 3 demonstrates that $\hat{q}_{\pi}$ corresponds to the state-action value function of the policy $\pi$ but for a different reward function $\hat{r}$ (c.f., (7)). Thus, designing a lower-variance behaviour policy boils down to learning an additional Q-value network to estimate $\hat{q}_{\pi}$ (also $q_{\pi}$ – if not already available) and optimizing the behaviour policy $\mu$ so that it matches (5).
@@ 5.2 Fitted Q-evaluation
The stability of the estimates for $q_{\pi}$ and $\hat{q}_{\pi}$ are critical to BPO. Value-based approaches, including FQE, are known to be prone to overestimation and approximation bias (fujimoto2022should). To stabilize learning we use symlog targets (hafner2024mastering), where the targets $Q^{\text{targ}}_{t}$ and $\hat{Q}^{\text{targ}}_{t}$ are transformed with the symmetric log: $\text{symlog}(x)=\text{sign}(x)\ln(|x|+1)$. To preserve their magnitudes the predictions are transformed back with the symmetric exp: $\text{symexp}(x)=\text{sign}(x)(\exp(|x|)-1)$. Symlog targets squash large magnitudes allowing the two networks to quickly move their predictions to large values when needed. This is crucial as the relative magnitude between $Q_{\zeta}$ and $\hat{Q}_{\hat{\zeta}}$ predictions can be large and symlog targets allow us to use the same hyperparameters (e.g., learning rate) for both $Q_{\zeta}$ and $\hat{Q}_{\hat{\zeta}}$ and keep the learning dynamics consistent for both networks. Further techniques used to deal with overestimation and stability issues are detailed in Appendix C.2.
@@ 5.3 Behaviour Policy
We establish the following result for this loss function, the proof can be found in Appendix A.8
@@ 6 Experimental Results
In this section we present our experimental results. We start with REINFORCE (williams1992simple), which is a simple algorithm that uses the full Monte Carlo returns (i.e., $\lambda=1$) to update the policy with the policy gradient (c.f., (9)). Given the relatively simplicity of REINFORCE we test our approach on a simple, yet non-trivial toy example from (sutton2018reinforcement): the ShortCorridor grid world. Given the example is simple, the gains here are marginal, and these results are simply to demonstrate that we can use our method for a variety of policy-gradient algorithms. We then continue by presenting our results for BPO on much more complex environments, where the underlying algorithm we build on top of PPO (schulman2017proximalpolicyoptimizationalgorithms) – a more sophisticated and widely used algorithm. Additional details of both the REINFORCE and PPO update are in Appendix C.1.
@@
We now evaluate BPO on several MuJoCo (todorov2012mujoco) environments provided by Gymnasium (towers2025gymnasium), testing also, different hyperparameter settings for the underlying PPO algorithm. In particular we test with the standard hyperparameter settings (default) provided in (schulman2017proximalpolicyoptimizationalgorithms), we also test with generalized state-dependent exploration (gSDE) (raffin2022smooth). All hyperparameter settings are detailed in Appendix C.2, we note that for a fair comparison between BPO and PPO, we keep the underlying PPO parameters identical and we only tune the BPO specific hyperparameters (e.g., $\bar{c}$, $\bar{\rho}$), although in practice one might be able to push the PPO hyperparameters further (e.g., $\lambda\to 1$) as BPO reduces variance. The results are summarized in Table 1. We also provide complete learning curves for each environment, see Figure 3.
@@ Discussion.
In our experiments we search over various hyperparameters settings for the truncation parameters $\bar{c}$ and $\bar{\rho}$, we also experiment with trajectory level truncation (traj), where the full product $(\prod^{k-1}_{i=t}c_{i})\rho_{k}$ in (6) is truncated with $\bar{c}$, rather than truncating each individual IS ratio. The advantage of this, is ratios greater than $\bar{c}$ can propagate through the trajectory, but the product is always truncated for stable learning. Again, this can be efficiently computing by stepping backward through time. In Table 1 we see that almost all configurations of BPO improve upon the baseline (PPO), and in most cases these improvements are statistically significant. Studying Figure 3 we see that BPO often exhibits faster convergence at the start of training and is more stable towards the end, demonstrating the well-designed behaviour policies can not only theoretically but empirically improve the convergence and stability of online RL, by reducing the variance of the returns estimator. We conduct some ablations for BPO, to observe the effect of removing the $\hat{q}_{\pi}$ estimate from the loss function in (16) and investigating the effect of symlog targets. For these results please see Appendix B.
@@ 7 Conclusion
In this paper we demonstrate that well designed behaviour policies can provably reduce the variance of practical returns estimators for online RL. We develop a sound methodology for training variance-reducing behaviour policies which are used to collect off-policy data for concurrently optimizing a target policy. Empirically, we show that our approach enables stable and more efficient off-policy RL, compared to the on-policy counterparts. Directions for future work include, extending our methodology to other policy-gradient algorithms, e.g., A2C (mnih2016asynchronous). It would also be worth exploring if our methodology is useful for value-based methods, e.g., DDPG (silver2014deterministic), TD3 (fujimoto2018addressing) and SAC (haarnoja2018soft). Furthermore, an ambitious goal might be to use model-based RL (sutton1990integrated) to assist in designing globally optimal behaviour policies.

### v2

@@ 3.3 Variance reduction
Figure 1: (a) Sampling densities. (b) Distribution of estimates over 100,000 trials.
@@ 4 Off-policy TD($\lambda$)
This property is important, as with appropriate hyperparameter settings we can achieve unbiasedness. We treat $\bar{c}$ and $\bar{\rho}$ as hyperparameters; generally this result indicates that a good approach is to set $\bar{\rho}\geq\bar{c}\geq 1$. Additionally, $\lambda$ should not stray too far from 1 as it biases our estimate when we rely on value estimates. We continue with the the following result.
@@
Theorem 1 ensures that while we search over the space of policies $\Lambda$ our estimator remains unbiased (under certain conditions). The full proof can be found in the extended version of the paper. Now we provide provable variance reduction for $G_{t}^{\text{TIS}}$ under the behaviour policy $\hat{\mu}$ designed in (5).
@@
And, $\epsilon(s):=c(s)+\gamma^{2}\mathbb{E}_{A_{t}\sim\hat{\mu}_{t}}\Big[\rho_{t}^{2}\mathbb{E}_{S_{t+1}\sim p}\big[\epsilon(S_{t+1})\big]\Big]$
The full proof can be found in the extended version of the paper. Noting here that $c(s)$ is always non-negative by Jensen’s inequality, which guarantees non-negativity of $\epsilon(s)$ and thus the property of variance reduction. $c(s)=0$ occurs only in the degenerate case when all actions have the same $\hat{q}_{\pi}$ for state $s$. It remains to show how $\hat{q}_{\pi}$ can be learnt,
@@
The proof here is a straightforward adaptation of (liu2024efficient) and intermediate results from the other theorems (see extended version). Theorem 3 demonstrates that $\hat{q}_{\pi}$ corresponds to the state-action value function of the policy $\pi$ but for a different reward function $\hat{r}$ (c.f., (7)). Thus, designing a lower-variance behaviour policy boils down to learning an additional Q-value network to estimate $\hat{q}_{\pi}$ (also $q_{\pi}$ – if not already available) and optimizing the behaviour policy $\mu$ so that it matches (5).
@@ 5.2 Fitted Q-evaluation
The stability of the estimates for $q_{\pi}$ and $\hat{q}_{\pi}$ are critical to BPO. Value-based approaches, including FQE, are known to be prone to overestimation and approximation bias (fujimoto2022should). To stabilize learning we use symlog targets (hafner2024mastering), where the targets $Q^{\text{targ}}_{t}$ and $\hat{Q}^{\text{targ}}_{t}$ are transformed with the symmetric log: $\text{symlog}(x)=\text{sign}(x)\ln(|x|+1)$. To preserve their magnitudes the predictions are transformed back with the symmetric exp: $\text{symexp}(x)=\text{sign}(x)(\exp(|x|)-1)$. Symlog targets squash large magnitudes allowing the two networks to quickly move their predictions to large values when needed. This is crucial as the relative magnitude between $Q_{\zeta}$ and $\hat{Q}_{\hat{\zeta}}$ predictions can be large and symlog targets allow us to use the same hyperparameters (e.g., learning rate) for both $Q_{\zeta}$ and $\hat{Q}_{\hat{\zeta}}$ and keep the learning dynamics consistent for both networks. Further techniques used to deal with overestimation and stability issues can be found in the appendix of the extended version of the paper.
@@ 5.3 Behaviour Policy
We establish the following result for this loss function, the proof can be found in the extended version.
@@ 6 Experimental Results
In this section we present our experimental results. We start with REINFORCE (williams1992simple), which is a simple algorithm that uses the full Monte Carlo returns (i.e., $\lambda=1$) to update the policy with the policy gradient (c.f., (9)). Given the relatively simplicity of REINFORCE we test our approach on a simple, yet non-trivial toy example from (sutton2018reinforcement): the ShortCorridor grid world. Given the example is simple, the gains here are marginal, and these results are simply to demonstrate that we can use our method for a variety of policy-gradient algorithms. We then continue by presenting our results for BPO on much more complex environments, where the underlying algorithm we build on top of PPO (schulman2017proximalpolicyoptimizationalgorithms) – a more sophisticated and widely used algorithm.
@@
We now evaluate BPO on several MuJoCo (todorov2012mujoco) environments provided by Gymnasium (towers2025gymnasium), testing also, different hyperparameter settings for the underlying PPO algorithm. In particular we test with the standard hyperparameter settings (default) provided in (schulman2017proximalpolicyoptimizationalgorithms), we also test with generalized state-dependent exploration (gSDE) (raffin2022smooth). All hyperparameter settings are detailed in the appendix of the extended version, we note that for a fair comparison between BPO and PPO, we keep the underlying PPO parameters identical and we only tune the BPO specific hyperparameters (e.g., $\bar{c}$, $\bar{\rho}$), although in practice one might be able to push the PPO hyperparameters further (e.g., $\lambda\to 1$) as BPO reduces variance. The results are summarized in Table 1. We also provide complete learning curves for each environment, see Figure 3.
@@ Discussion.
In our experiments we search over various hyperparameters settings for the truncation parameters $\bar{c}$ and $\bar{\rho}$, we also experiment with trajectory level truncation (traj), where the full product $(\prod^{k-1}_{i=t}c_{i})\rho_{k}$ in (6) is truncated with $\bar{c}$, rather than truncating each individual IS ratio. The advantage of this, is ratios greater than $\bar{c}$ can propagate through the trajectory, but the product is always truncated for stable learning. Again, this can be efficiently computing by stepping backward through time. In Table 1 we see that almost all configurations of BPO improve upon the baseline (PPO), and in most cases these improvements are statistically significant. Studying Figure 3 we see that BPO often exhibits faster convergence at the start of training and is more stable towards the end, demonstrating the well-designed behaviour policies can not only theoretically but empirically improve the convergence and stability of online RL, by reducing the variance of the returns estimator. In the extended version of the paper we conduct some ablations for BPO, to observe the effect of removing the $\hat{q}_{\pi}$ estimate from the loss function in (16) and investigating the effect of symlog targets.
@@ 7 Conclusion
In this paper we demonstrate that well designed behaviour policies can provably reduce the variance of practical returns estimators for online RL. We develop a sound methodology for training variance-reducing behaviour policies which are used to collect off-policy data for concurrently optimizing a target policy. Empirically, we show that our approach enables stable and more efficient off-policy RL, compared to the on-policy counterparts. Directions for future work include, extending our methodology to other policy-gradient algorithms, e.g., A2C (mnih2016asynchronous). It would also be worth exploring if our methodology is useful for value-based methods, e.g., DDPG (silver2014deterministic), TD3 (fujimoto2018addressing) and SAC (haarnoja2018soft). We might also use model-based RL (sutton1990integrated) to assist in designing globally optimal behaviour policies.